the age of a father is now 5 times the age of his son. After 18 years, the age of the father will be twice his son's age.
a). Let x and y be the present age of the father and his son respectively. Set up two equations connecting x and y.
b). Find their present ages.

Answer by mananth(16946)

x=5y
x-5y=0..................(1)
after 18 years
father's age = (x+18)
son's age (y+18)
(x+18)=2(y+18)
x+18=2y+36
x-2y=18...............(2)
subtract (2) from (1)
-3y=-18
y=6
father = 5x = 5*6=30 years